OpenStudy (anonymous):

Will give medal! The domain of the following function is {-2, 0, 1.5, }. Find the range. Be sure to show all your work. f(x) = 4x – 3

OpenStudy (anonymous):

f(X)=4(-2)-3=-11 f(X)=4(0)-3=-3 f(X)=4(1.5)-3=3 Therefore, the range is (-11, -3, 3) in respect to the domain